OpenStudy (anonymous):

write the expression in terms of sin theta and cos theta. 1-sin squared(-theta) divided by 1+ cot squared (-theta)

OpenStudy (gorv):

\[\frac{ 1-\sin^{2}(-\theta) }{ 1+\cot^{2}(-\theta) }\]

OpenStudy (gorv):

sin^2(-theta)=sin^2Theta same for cot

OpenStudy (gorv):

\[\frac{ 1-\sin^2 \theta }{ 1+\cot^2\theta }\]

OpenStudy (gorv):

cosec^2(theta)=1+cot^2 (theta) and cosec^2=1/sin^2

OpenStudy (gorv):

1-sin^2(theta)=cos^2(theta)

OpenStudy (gorv):

\[\frac{ \cos^2\theta }{ \frac{ 1 }{ \sin^2\theta } }=\sin^2\theta*\cos^2\theta\]